A basket contains 10 ripe and 4 unripe bananas. If three bananas are taken from the basket one after the other, determine the possible values of the random variable U representing the number of unripe bananas


1
Expert's answer
2022-01-30T16:22:51-0500

Let "R" represent the ripe bananas and let "U" represent the unripe bananas.

List the sample space of this experiment


"S=\{RRR, RRU, RUR, URR,"

"RUU, URU, UUR, UUU\}"

The values of random variable "U" are "0, 1, 2, 3."

There are 8 possible outcomes:

one with "U=0\ (RRR),"

three with "U=1\ (RRU, RUR, URR),"

three with "U=2\ (RUU, URU, UUR),"

one with "U=3\ (UUU)."

Then


"P(U=0)=\dfrac{1}{8}"

"P(U=1)=\dfrac{3}{8}"

"P(U=2)=\dfrac{3}{8}"

"P(U=3)=\dfrac{1}{8}"

Construct the probability distribution of random variable "U"


"\begin{matrix}
 u & 0 & 1 & 2 & 3 \\
 p(u) & 1/8 & 3/8 & 3/8 & 1/8
\end{matrix}"